# How to Train Long-Context Language Models (Effectively)

### v3

@@ 1 Introduction
While many works have focused on extending the context length of pre-trained LMs with minimal training (Chen et al., 2023; Peng et al., 2024, inter alia), the above methods fail to solve even the simple needle-in-a-haystack (NIAH; Kamradt, 2024) task and it is necessary to continue LLM training on billions of tokens of long documents to learn this task robustly. Frontier open-source models, such as Llama-3.1 (Dubey et al., 2024) and Jamba (Lenz et al., 2024), also employ a long-context continued training stage, followed by supervised fine-tuning (SFT) on instruction data. We adopt the same setting and study continued training and SFT of a pre-trained LM for effective long-context use.
@@
Guided by our evaluation protocol, we run comprehensive experiments with Llama-3-8B (8K original context window; Dubey et al., 2024) to study each component of long-context continued training, including data mixture, data and length scaling, supervised fine-tuning, and many other design choices such as cross-document attention masking and position extrapolation. Many of our findings are surprising or contradictory to existing claims, for example, (1) training only on long data hurts long-context performance, (2) training on longer sequences than the evaluation length helps, and (3) SFT on only short instruction data is sufficient for good long-context performance. We outline our main takeaways and the structure of the paper in the takeaway box at the beginning of this section.
Our final model, ProLong, achieves the best performance at a 128K context length among 10B-parameter models, while taking only $5\%$ of the data budget compared to Llama-3.1’s long-context training (Dubey et al., 2024). ProLong has a maximum context length of 512K tokens, making it one of the longest-context LMs available. [^2] Throughout the paper, we use binary prefixes K$=2^{10}$, M=$2^{20}$, and B=$2^{30}$.
@@ 2 Guiding Model Development With Meaningful Evaluations
A pre-requisite for training a strong LM is having a robust evaluation suite that can guide model development while tracking its utility in real-world applications. While synthetic benchmarks like needle-in-a-haystack (NIAH; Kamradt, 2024) and RULER (Hsieh et al., 2024) have gained much popularity due to their simplicity and controllability, we are interested in a wider range of tasks that reflect practical usage, such as the ability to reason over the whole document. In the following, we describe our evaluation protocols and showcase why they are critical to our model development.
@@ 2.1 Evaluating on diverse and realistic tasks
We first make the decision to use HELMET (Yen et al., 2025) as our evaluation suite, as it is one of the most comprehensive long-context benchmarks. For fast iteration, we only use a subset of HELMET tasks for model development:
@@
We showcase the importance of a robust evaluation suite in Table 1. As a predecessor of our work, Fu et al. (2024) only consider needle-in-a-haystack (NIAH) and perplexity during model development; evaluations on 3 tasks from HELMET reveal major short-comings of their models despite perfect NIAH scores. We also see how NIAH and even the HELMET recall task become saturated for strong models (Llama-3.1-8B vs. 70B) while other task categories continue to detect differences in their long-context abilities.
Table 1: HELMET offers a more holistic long-context evaluation. We reproduce Fu et al. (2024) on Llama-3-8B with SFT. We report the instruct Llama versions.
@@
Besides synthetic recall tasks, many previous works rely on perplexity (PPL) for evaluating long-context extensions of LMs (Chen et al., 2023; Fu et al., 2024; Lu et al., 2024), which is commonly measured on the PG19 books dataset (Rae et al., 2020). We use the ablation experiment from §3.2 to showcase why perplexity is not an indicative metric for developing long-context models. The experiment studies how the ratio of long documents affects the performance. We report both our evaluation and the perplexity measured on the last 32K tokens of 64K-length documents from PG19. As shown in Figure 1, while using more long data continues to improve PPL, it is clear that using 100% long data significantly hurts downstream long-context performance.
@@ 2.2 Evaluating after supervised fine-tuning
Supervised fine-tuning (SFT; Ouyang et al., 2022) is an additional training stage that fine-tunes the model on a small amount of natural-language instructions and corresponding responses; it enables a base LM to address user queries in a chat format and has become a standard step for producing frontier LMs. Here, we consider the difference between evaluating a model before or after SFT.
In preliminary experiments, we continue training Llama-3-8B-Base on 5B-token subsets from the data mix by Fu et al. (2024). The mix is based on SlimPajama (Soboleva et al., 2023) and upsamples long documents to constitute roughly 70% of tokens, while retaining the original domain proportions. Then we conduct SFT on several intermediate checkpoints with UltraChat (Ding et al., 2023).
@@ 2.3 Checking that short-context performance is preserved
Long-context abilities should not come at the expense of short-context performance, particularly since short-context evaluations cover a wider range of capabilities, e.g., world knowledge, commonsense, and mathematical reasoning. However, short-context evaluation has largely been neglected by previous long-context research. We report on 5 tasks from the the Open LLM Leaderboard (Beeching et al., 2023): HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), ARC-challenge (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), and GSM8K (Cobbe et al., 2021). We evaluate short-context performance before SFT, since this allows for a direct comparison to the base model which was used as initialization for the long-context training.
@@
We show in Table 2 that both training-free position extrapolation, as well as fine-tuning with an existing long data mixture (Fu et al., 2024) do not preserve the strong performance of Llama-3-8B on standard short-context tasks. This motivates us to find data sources which retain the initial model’s strong short-context performance.
Table 2: Applying position extrapolation (PE) to Llama-3-8B by changing the RoPE frequency base (§B.1) or fine-tuning it on a long-context SlimPajama mixture (Fu et al., 2024; Soboleva et al., 2023) deteriorates the performance of this top-shelf pre-trained LM on short-context tasks.
@@ 3 Long-Context Data Curation
The quality and composition of training data has been found to be the most important factor for LM pre-training (Penedo et al., 2023; Wettig et al., 2024; Li et al., 2024a) and is therefore a primary focus of our study. To make data decisions, we perform ablation experiments: we continue to train Llama-3-8B-Base for 5B tokens with a maximum length of 64K tokens and evaluate according to §2. See §A.4 for more details of our ablation setting.
@@ 3.1 Code repositories and books are good sources of long-context data
We analyze the quantity of long data in SlimPajama (SP; Soboleva et al., 2023). Table 3 shows that books account for the majority of long-context tokens. When inspecting the long data in CommonCrawl (CC), we observe that though varied in quality, it also contains some book-like content, which future work could identify via data selection methods.
@@
While only few files from GitHub reach a very long length (which also tend to be lower quality as suggested by Singh et al., 2024), we construct an abundant source of long-context data from the Stack (Kocetkov et al., 2023) by concatenating all files from a repository to form a single document. Unlike Guo et al. (2024), we do not order the files based on dependencies, which should increase the distance between dependent files and reduce recency bias.
@@ 3.2 Training only on long data hurts long-context performance
The ratio between short/long data is another crucial factor for downstream performance. Prior work either trains only on long data (Peng et al., 2024) or adds some short training data (Yen et al., 2024; Fu et al., 2024). However, we are the first to systematically study the impact of short/long ratio.
@@ 3.3 Choosing a high-quality short-context mix is important
We saw in §2.3 that it is difficult to preserve the strong performance of Llama-3-8B on short-context tasks during long-context fine-tuning. We adopt our best long-context settings (Book/repo data and 60% long/40% short) and study the impact of different short-context training mixes. We experiment with SlimPajama (Soboleva et al., 2023), FineWeb-Edu (Penedo et al., 2024), DCLM-Baseline (Li et al., 2024a), and our own ProLong ShortMix. Our ShortMix is inspired by the “stage 2 training” in MiniCPM (Hu et al., 2024a) and Dolma-1.7 (Soldaini et al., 2024), which use more knowledge-intensive, downstream-related data at the end of pre-training. Table 5 shows the composition of our ShortMix.333 Since we do not truncate documents in the short data component unnecessarily, it includes a small percentage of documents longer than 8K. See Table 14 in the appendix for the dataset length statistics.
@@
To confirm the effectiveness of our long-context data curation, we conduct a head-to-head comparison with the previous work by Fu et al. (2024) in §B.6. Our data mix significantly surpasses Fu et al. (2024) in both long-context and short-context tasks, underscoring the efficacy of ProLong.
@@ 4 Scaling the Size and Length of the Training Data
One might assume that we should train long-context models on the maximum sequence length that we want the model to support. Many works even emphasize extrapolation to even longer sequences at inference time (Press et al., 2022; Xiao et al., 2024b; a; Yen et al., 2024; Chen et al., 2023). In contrast, we observe that training on a longer sequence length (512K tokens) substantially improves the long-context performance at a shorter evaluation length (64K tokens).
@@
What is the benefit of training on longer sequences than used during evaluation? Here is our hypothesis: Assume that in order to solve a certain task (e.g., recall), a model has to be trained on examples of dependencies that span a distance of precisely $d$ tokens, i.e., there is no generalization between dependencies of different lengths. Also assume that these dependencies are equally likely to occur at any position in a sequence. Then, a document of length $nd$ will have $(n-1)(d-1)$ more dependencies of distance $d$ than $n$ documents of length $d$. While these assumptions do not hold in practice, this simplified model still provides intuition for our empirical findings.
@@ 5 Supervised Fine-Tuning for Long-Context LMs
In this section, we study how to best enable long-context language models to follow instructions. We focus on supervised fine-tuning on instruction datasets (Ouyang et al., 2022) and leave reinforcement learning and preference optimization for future work.
All our experiments in this section use the ProLong base model, which was trained for 40B tokens at a maximum sequence length of 512K. In comparison, open-source instruction data are very short, e.g., UltraChat (Ding et al., 2023) conversations have 1.2K tokens on average and 4.1K tokens maximum. To bridge this gap, several works (Xiong et al., 2023; Dubey et al., 2024; Xiong et al., 2024) have proposed to generate long instruction data synthetically.
We consider three popular SFT datasets—UltraChat (Ding et al., 2023), Tulu-v2 (Ivison et al., 2023), ShareGPT [^5] https://huggingface.co/datasets/RyokoAI/ShareGPT52K.—and three sources of synthetic data: For synthetic QA, we prompt Llama-3-8B-Instruct to generate a question-and-answer pair given a random chunk from a long document; we reuse the QA pairs for synthetic RAG but we present a random list of chunks from the document to mimic retrieved passages; for synthetic summarization, we generate summaries for long books via recursive summarization (Wu et al., 2021). For all synthetic data, we write several templates, which we sample at random to increase diversity. More details can be found in §A.5. We always use a combination of 40% synthetic QA, 30% synthetic RAG, and 30% synthetic summarization in our synthetic instruction dataset. The hyperparameters for the instruction tuning experiments can be found in Table 9.
@@
Why do our conclusions about synthetic data differ from previous work? We offer the following hypotheses: (1) Previous work like Xiong et al. (2024); Bai et al. (2024a) may have insufficient long-context training and the synthetic data acts as additional long-context training data. (2) Our instruction dataset is much smaller compared to the private instruction data used for Llama-3.1 (Dubey et al., 2024)—it is possible that when using an extensive short instruction dataset, mixing in synthetic long data avoids the model from degenerating on long-context tasks.
@@
LR: $51e-$ 5 with $10\%$ warmup and cosine decay to $61e-$ 6, each stage
@@
LR = $52e-$ 5 (cosine decay to $62e-$ 6), warmup = $5\%$
@@ 6.1 Final recipe
We summarize the training recipe for ProLong in Table 9. Our final model starts from the Llama-3-8B-Instruct model and is trained on 64K sequence length for 20B tokens. It is then further trained on 512K sequence length for 20B tokens (ProLong base), which we achieve using sequence parallelism (Li et al., 2023). We obtain the final ProLong model via SFT of the base model on UltraChat. One small difference on the data mixture between our ablations and the final model is that we mix in 3% high-quality textbooks (Chevalier et al., 2024), as book-like data are shown to be beneficial for long-context (§3.1) and textbooks are highly educational. This also slightly changes the proportions of ShortMix. You can find more details about our data processing (§A.2) and the training stack (§A.3) in the appendix.
@@
We find that changing the RoPE (Su et al., 2021) frequency base to achieve position extrapolation (Xiong et al., 2023; emozilla, 2023) significantly improves long-context performance, even with a significant amount of training. §B.1 shows our ablation on the best RoPE base to use. While the original Llama models use a RoPE base of $10^{5}$, we use a base of $8\times 10^{6}$ for the 64K setting and $1.28\times 10^{8}$ for the 512K setting.
@@
Ding et al. (2024a) show that masking out attention across document boundaries improve model performance and this was also used during Llama-3 pre-training (Dubey et al., 2024). In §B.2, we show that disabling cross-document attention in continued training benefits both the short and long-context performance. Disabling cross-document attention can also result in higher training throughput, which we describe in more detail in §A.3.

### v4

@@ 1 Introduction
While many works have focused on extending the context length of pre-trained LMs with minimal training (chen2023extending; peng2024yarn, inter alia), the above methods fail to solve even the simple needle-in-a-haystack (NIAH; gkamradt_llmtest_needleinahaystack_2024) task and it is necessary to continue LLM training on billions of tokens of long documents to learn this task robustly. Frontier open-source models, such as Llama-3.1 (dubey2024llama) and Jamba (team2024jamba15), also employ a long-context continued training stage, followed by supervised fine-tuning (SFT) on instruction data. We adopt the same setting and study continued training and SFT of a pre-trained LM for effective long-context use.
@@
Guided by our evaluation protocol, we run comprehensive experiments with Llama-3-8B (8K original context window; dubey2024llama) to study each component of long-context continued training, including data mixture, data and length scaling, supervised fine-tuning, and many other design choices such as cross-document attention masking and position extrapolation. Many of our findings are surprising or contradictory to existing claims, for example, (1) training only on long data hurts long-context performance, (2) training on longer sequences than the evaluation length helps, and (3) SFT on only short instruction data is sufficient for good long-context performance. We outline our main takeaways and the structure of the paper in the takeaway box at the beginning of this section.
[…]
@@ 2 Guiding Model Development With Meaningful Evaluations
A pre-requisite for training a strong LM is having a robust evaluation suite that can guide model development while tracking its utility in real-world applications. While synthetic benchmarks like needle-in-a-haystack (NIAH; gkamradt_llmtest_needleinahaystack_2024) and RULER (hsieh2024ruler) have gained much popularity due to their simplicity and controllability, we are interested in a wider range of tasks that reflect practical usage, such as the ability to reason over the whole document. In the following, we describe our evaluation protocols and showcase why they are critical to our model development.
@@ 2.1 Evaluating on diverse and realistic tasks
We first make the decision to use HELMET (yen2024helmet) as our evaluation suite, as it is one of the most comprehensive long-context benchmarks. For fast iteration, we only use a subset of HELMET tasks for model development:
@@
We showcase the importance of a robust evaluation suite in Table 1. As a predecessor of our work, fu2024data only consider needle-in-a-haystack (NIAH) and perplexity during model development; evaluations on 3 tasks from HELMET reveal major short-comings of their models despite perfect NIAH scores. We also see how NIAH and even the HELMET recall task become saturated for strong models (Llama-3.1-8B vs. 70B) while other task categories continue to detect differences in their long-context abilities.
Table 1: HELMET offers a more holistic long-context evaluation. We reproduce fu2024data on Llama-3-8B with SFT. We report the instruct Llama versions.
@@
Besides synthetic recall tasks, many previous works rely on perplexity (PPL) for evaluating long-context extensions of LMs (chen2023extending; fu2024data; lu2024controlled), which is commonly measured on the PG19 books dataset (Rae2020Compressive). We use the ablation experiment from §3.2 to showcase why perplexity is not an indicative metric for developing long-context models. The experiment studies how the ratio of long documents affects the performance. We report both our evaluation and the perplexity measured on the last 32K tokens of 64K-length documents from PG19. As shown in Figure 1, while using more long data continues to improve PPL, it is clear that using 100% long data significantly hurts downstream long-context performance.
@@ 2.2 Evaluating after supervised fine-tuning
Supervised fine-tuning (SFT; ouyang2022training) is an additional training stage that fine-tunes the model on a small amount of natural-language instructions and corresponding responses; it enables a base LM to address user queries in a chat format and has become a standard step for producing frontier LMs. Here, we consider the difference between evaluating a model before or after SFT.
In preliminary experiments, we continue training Llama-3-8B-Base on 5B-token subsets from the data mix by fu2024data. The mix is based on SlimPajama (cerebras2023slimpajama) and upsamples long documents to constitute roughly 70% of tokens, while retaining the original domain proportions. Then we conduct SFT on several intermediate checkpoints with UltraChat (ding-etal-2023-enhancing).
@@ 2.3 Checking that short-context performance is preserved
Long-context abilities should not come at the expense of short-context performance, particularly since short-context evaluations cover a wider range of capabilities, e.g., world knowledge, commonsense, and mathematical reasoning. However, short-context evaluation has largely been neglected by previous long-context research. We report on 5 tasks from the the Open LLM Leaderboard (open-llm-leaderboard): HellaSwag (zellers2019hellaswag), MMLU (hendrycks2021measuring), ARC-challenge (clark2018think), WinoGrande (sakaguchi2021winogrande), and GSM8K (cobbe2021training). We evaluate short-context performance before SFT, since this allows for a direct comparison to the base model which was used as initialization for the long-context training.
@@
We show in Table 2 that both training-free position extrapolation, as well as fine-tuning with an existing long data mixture (fu2024data) do not preserve the strong performance of Llama-3-8B on standard short-context tasks. This motivates us to find data sources which retain the initial model’s strong short-context performance.
Table 2: Applying position extrapolation (PE) to Llama-3-8B by changing the RoPE frequency base (§B.1) or fine-tuning it on a long-context SlimPajama mixture (fu2024data; cerebras2023slimpajama) deteriorates the performance of this top-shelf pre-trained LM on short-context tasks.
@@ 3 Long-Context Data Curation
The quality and composition of training data has been found to be the most important factor for LM pre-training (penedo2023refinedweb; wettig2024qurating; li2024datacomplm) and is therefore a primary focus of our study. To make data decisions, we perform ablation experiments: we continue to train Llama-3-8B-Base for 5B tokens with a maximum length of 64K tokens and evaluate according to §2. See §A.4 for more details of our ablation setting.
@@ 3.1 Code repositories and books are good sources of long-context data
We analyze the quantity of long data in SlimPajama (SP; cerebras2023slimpajama). Table 3 shows that books account for the majority of long-context tokens. When inspecting the long data in CommonCrawl (CC), we observe that though varied in quality, it also contains some book-like content, which future work could identify via data selection methods.
@@
While only few files from GitHub reach a very long length (which also tend to be lower quality as suggested by singh2024brevity), we construct an abundant source of long-context data from the Stack (kocetkov2023the) by concatenating all files from a repository to form a single document. Unlike guo2024deepseek, we do not order the files based on dependencies, which should increase the distance between dependent files and reduce recency bias.
@@ 3.2 Training only on long data hurts long-context performance
The ratio between short/long data is another crucial factor for downstream performance. Prior work either trains only on long data (peng2024yarn) or adds some short training data (yen-etal-2024-long; fu2024data). However, we are the first to systematically study the impact of short/long ratio.
@@ 3.3 Choosing a high-quality short-context mix is important
We saw in §2.3 that it is difficult to preserve the strong performance of Llama-3-8B on short-context tasks during long-context fine-tuning. We adopt our best long-context settings (Book/repo data and 60% long/40% short) and study the impact of different short-context training mixes. We experiment with SlimPajama (cerebras2023slimpajama), FineWeb-Edu (penedo2024finewebdatasetsdecantingweb), DCLM-Baseline (li2024datacomplm), and our own ProLong ShortMix. Our ShortMix is inspired by the “stage 2 training” in MiniCPM (hu2024minicpm) and Dolma-1.7 (soldaini-etal-2024-dolma), which use more knowledge-intensive, downstream-related data at the end of pre-training. Table 5 shows the composition of our ShortMix. [^3] Since we do not truncate documents in the short data component unnecessarily, it includes a small percentage of documents longer than 8K. See Table 14 in the appendix for the dataset length statistics.
@@
To confirm the effectiveness of our long-context data curation, we conduct a head-to-head comparison with the previous work by fu2024data in §B.6. Our data mix significantly surpasses fu2024data in both long-context and short-context tasks, underscoring the efficacy of ProLong.
@@ 4 Scaling the Size and Length of the Training Data
One might assume that we should train long-context models on the maximum sequence length that we want the model to support. Many works even emphasize extrapolation to even longer sequences at inference time (press2022train; xiao2024efficient; xiao2024infllm; yen-etal-2024-long; chen2023extending). In contrast, we observe that training on a longer sequence length (512K tokens) substantially improves the long-context performance at a shorter evaluation length (64K tokens).
@@
What is the benefit of training on longer sequences than used during evaluation? Here is our hypothesis: Assume that in order to solve a certain task (e.g., recall), a model has to be trained on examples of dependencies that span a distance of precisely $d$ tokens, i.e., there is no generalization between dependencies of different lengths. Also assume that these dependencies are equally likely to occur at any position in a sequence. Then, a document of length $nd$ will have (n1)(d1)(n-1)(d-1) more dependencies of distance $d$ than $n$ documents of length $d$. While these assumptions do not hold in practice, this simplified model still provides intuition for our empirical findings.
@@ 5 Supervised Fine-Tuning for Long-Context LMs
In this section, we study how to best enable long-context language models to follow instructions. We focus on supervised fine-tuning on instruction datasets (ouyang2022training) and leave reinforcement learning and preference optimization for future work.
All our experiments in this section use the ProLong base model, which was trained for 40B tokens at a maximum sequence length of 512K. In comparison, open-source instruction data are very short, e.g., UltraChat (ding-etal-2023-enhancing) conversations have 1.2K tokens on average and 4.1K tokens maximum. To bridge this gap, several works (xiong2023effective; dubey2024llama; xiong2024artificial) have proposed to generate long instruction data synthetically.
We consider three popular SFT datasets—UltraChat (ding-etal-2023-enhancing), Tulu-v2 (ivison2023camels), ShareGPT [^5] https://huggingface.co/datasets/RyokoAI/ShareGPT52K.—and three sources of synthetic data: For synthetic QA, we prompt Llama-3-8B-Instruct to generate a question-and-answer pair given a random chunk from a long document; we reuse the QA pairs for synthetic RAG but we present a random list of chunks from the document to mimic retrieved passages; for synthetic summarization, we generate summaries for long books via recursive summarization (wu2021recursively). For all synthetic data, we write several templates, which we sample at random to increase diversity. More details can be found in §A.5. We always use a combination of 40% synthetic QA, 30% synthetic RAG, and 30% synthetic summarization in our synthetic instruction dataset. The hyperparameters for the instruction tuning experiments can be found in Table 9.
@@
Why do our conclusions about synthetic data differ from previous work? We offer the following hypotheses: (1) Previous work like xiong2024artificial; bai2024longalign may have insufficient long-context training and the synthetic data acts as additional long-context training data. (2) Our instruction dataset is much smaller compared to the private instruction data used for Llama-3.1 (dubey2024llama)—it is possible that when using an extensive short instruction dataset, mixing in synthetic long data avoids the model from degenerating on long-context tasks.
@@
LR: 1e51e-5 with $10\%$ warmup and cosine decay to 1e61e-6, each stage
@@
LR = 2e52e-5 (cosine decay to 2e62e-6), warmup = $5\%$
@@ 6.1 Final recipe
We summarize the training recipe for ProLong in Table 9. Our final model starts from the Llama-3-8B-Instruct model and is trained on 64K sequence length for 20B tokens. It is then further trained on 512K sequence length for 20B tokens (ProLong base), which we achieve using sequence parallelism (li-etal-2023-sequence). We obtain the final ProLong model via SFT of the base model on UltraChat. One small difference on the data mixture between our ablations and the final model is that we mix in 3% high-quality textbooks (chevalier2024language), as book-like data are shown to be beneficial for long-context (§3.1) and textbooks are highly educational. This also slightly changes the proportions of ShortMix. You can find more details about our data processing (§A.2) and the training stack (§A.3) in the appendix.
@@
We find that changing the RoPE (su2021roformer) frequency base to achieve position extrapolation (xiong2023effective; dynamicntk) significantly improves long-context performance, even with a significant amount of training. §B.1 shows our ablation on the best RoPE base to use. While the original Llama models use a RoPE base of $10^{5}$, we use a base of $8\times 10^{6}$ for the 64K setting and $1.28\times 10^{8}$ for the 512K setting.
@@
ding2024fewer show that masking out attention across document boundaries improve model performance and this was also used during Llama-3 pre-training (dubey2024llama). In §B.2, we show that disabling cross-document attention in continued training benefits both the short and long-context performance. Disabling cross-document attention can also result in higher training throughput, which we describe in more detail in §A.3.